# PromptReps: Prompting Large Language Models to Generate Dense and Sparse Representations for Zero-Shot Document Retrieval

### v2

@@ Abstract
Utilizing large language models (LLMs) for zero-shot document ranking is done in one of two ways: 1) prompt-based re-ranking methods, which require no further training but are only feasible for re-ranking a handful of candidate documents due to computational costs; and 2) unsupervised contrastive trained dense retrieval methods, which can retrieve relevant documents from the entire corpus but require a large amount of paired text data for contrastive training. In this paper, we propose PromptReps, which combines the advantages of both categories: no need for training and the ability to retrieve from the whole corpus. Our method only requires prompts to guide an LLM to generate query and document representations for effective document retrieval. Specifically, we prompt the LLMs to represent a given text using a single word, and then use the last token’s hidden states and the corresponding logits associated with the prediction of the next token to construct a hybrid document retrieval system. The retrieval system harnesses both dense text embedding and sparse bag-of-words representations given by the LLM. We further explore variations of this core idea that consider the generation of multiple words, and representations that rely on multiple embeddings and sparse distributions. Our experimental evaluation on the MSMARCO, TREC deep learning and BEIR zero-shot document retrieval datasets illustrates that this simple prompt-based LLM retrieval method can achieve a similar or higher retrieval effectiveness than state-of-the-art LLM embedding methods that are trained with large amounts of unsupervised data, especially when using a larger LLM. [^1] Code for fully reproducing the results is available at https://github.com/ielab/PromptReps.
@@
Shengyao Zhuang1,3, Xueguang Ma2, Bevan Koopman1,3, Jimmy Lin2, Guido Zuccon3 1CSIRO, 2 University of Waterloo, 3 The University of Queensland
@@ 1 Introduction
Figure 1: Overview of PromptReps. LLMs are prompted to simultaneously generate dense and sparse representations, which are then used to build search indices.
One line of work focuses on directly prompting LLMs to infer document relevance to a given query Sachan et al. (2022); Zhuang et al. (2023a); Ma et al. (2023b); Sun et al. (2023); Pradeep et al. (2023); Zhuang et al. (2023b); Qin et al. (2024). For instance, RankGPT Sun et al. (2023) casts document re-ranking as a permutation generation task, prompting LLMs to generate re-ordered document identifiers according to the document’s relevance to the query. These methods leverage LLMs for document ranking in a complete zero-shot setting where no further training is required. However, these methods can only serve as a second-stage re-ranker on a handful of candidate documents. This is because each prompt requires one full LLM inference: for example, in the case of a corpus with 1M documents, a pointwise approach would require the construction of 1M prompts and thus the execution of 1M (costly) LLM inferences – making it unfeasible for an online search engine.
@@ 2.1 Supervised Neural Retrievers
Neural retrievers based on the bi-encoder architecture bring significant improvements over traditional best-match retrievers such as BM25. Dense retrievers such as DPR Karpukhin et al. (2020), ANCE Xiong et al. (2021), ColBERT Khattab and Zaharia (2020), are based on encoder-only language models and encode text into low-dimensional dense vectors, conducting search with (approximate) nearest neighbor search. On the other hand, sparse neural retrievers such as DeepImpact Mallia et al. (2021), uniCOIL Lin and Ma (2021), TILDE Zhuang and Zuccon (2021c, b), and SPLADE Formal et al. (2021), also based on encoder-only language models, encode text into high-dimensional sparse vectors as bag-of-words representations, conducting search in an inverted index. Recent work has also explored fine-tuning generative LLMs as dense retrievers such as RepLLaMA Ma et al. (2023a) and LLaRA Liao et al. (2024). A hybrid neural retrieval system refers to a system that combines the rankings provided by both dense and sparse retrievers, often resulting in an enhanced final ranking Lin and Ma (2021); Wang et al. (2021a).
@@
All these retrievers are trained with supervised relevance judgment data (e.g., MS MARCO Bajaj et al. (2018)) using contrastive learning. Our work instead focuses on building a hybrid neural retrieval system with zero-shot dense and sparse document representations without supervised contrastive learning and based on generative LLMs. This capability has two implications: (1) no contrastive training is required, which is expensive when applied to LLMs with several billions parameters, and (2) no human-labelled training data is required, which may be laborious and expensive to obtain. With regards to the first point, Wang et al. (2024b) reported that the training of E5-mistrail (7B parameters) took about 18 hours on 32 V100 GPUs, for an approximate cost of USD \$2,300222Based on 4 On-Demand p3dn.24xlarge instances, June 2024., emissions of $\approx$5.6 kgCO2e and consumption of $\approx$37.7 L of water for the associated cooling activities [^3] Emissions and water consumption estimates obtained using the frameworks of Scells et al. (2022); Zuccon et al. (2023).. Scaling this training to more and larger LLMs, and more data, will consequently further increase costs. Our proposed method does not incurr these additional contrastive pre-training costs. With regards to the second point, dense retrievers have shown to have poor generalisability when applied to data out-of-domain or out-of-task compared to the data used for contrastive training Thakur et al. (2021); Zhuang and Zuccon (2021a, 2022); Ren et al. (2023); Lin et al. (2023); Lupart et al. (2022). In presence of shift in data between training and deployment, retrieval losses can be significant: dense retrieval effectiveness can plummet far below that of best-match models like BM25 Khramtsova et al. (2023, 2024). The acquisition of in-domain/in-task training data can be costly, laborious and often impractical/impossible especially in domain-specific applications and when dealing with sensitive, private data.
@@ 2.2 Unsupervised Neural Retrievers
There have also been attempts at training effective neural retrievers without relying on human relevance judgments. Methods such as Contriever Izacard et al. (2022) and E5 Wang et al. (2024a), train a dense retriever with large-scale pseudo query-document pairs to build unsupervised (synthetic) training data. LLMs have also been adapted as unsupervised text embedding models for first-stage document retrieval. For instance, HyDE Gao et al. (2023a) enhances query representations for an unsupervised retriever by replacing the original query with LLM-generated hypothetical documents. More recent work has focused on directly converting generative LLMs into a text-embedding model with unsupervised contrastive pre-training. Methods like E5-Mistral-Inst Wang et al. (2024b) and Gecko Lee et al. (2024) use large-scale weakly supervised paired text data or LLM-generated query-document pair data to perform contrastive training on top of LLMs. LLM2Vec BehnamGhader et al. (2024), on the other hand, conducts further masked next token prediction pre-training with bidirectional attention, and SimCSE Gao et al. (2021) trains on raw text data to transform LLMs into text encoders. Although no labeled data is used, these methods require synthetic or unsupervised paired text data to perform contrastive pre-training (thus still experiencing training costs in terms of computations; and further computational costs may be associated with the generation of synthetic training data). Our method instead relies solely on prompt engineering to transform LLM into a robust text encoder for document retrieval without any extra training.
@@ 2.3 Prompting LLMs for document ranking
Inspired by the prompt-following capacity of LLMs, recent studies have explored prompting LLMs for document re-ranking. For instance, UPR Sachan et al. (2022) ranks documents pointwise by prompting the LLM to generate a relevant query for a given document and rank documents based on the likelihood of generating the query. RankGPT Sun et al. (2023) and LRL Ma et al. (2023b) propose to re-rank a list of documents at once and generate permutations for the reordered list. Pairwise Qin et al. (2024) and Setwise Zhuang et al. (2023b) prompting methods have also been explored to improve effectiveness and efficiency in the LLM re-ranking pipeline. These methods are only feasible for re-ranking a handful of candidate documents, thus limited to second-stage document re-ranking. In contrast, our approach utilizes prompts to construct the first-stage retrievers.
2.4 Prompting LLM for sentence embeddings
@@ 3 PromptReps
Table 1: nDCG@10 scores of BEIR 13 publicly available datasets. The best scores of methods without interpolating with BM25 are highlighted in bold.
Unsup Contrastive training PromptReps (ours)
LLM - BERT-330M Llama3-8B-I Llama3-8B-I Llama3-70B-I
Dataset BM25 E5-PTlarge LLM2Vec Dense Sparse Hybrid Dense Sparse Hybrid +BM25
arguana 39.70 44.4 51.73 29.70 22.85 32.98 31.65 24.66 35.27 39.53
climatefever 16.51 15.7 23.58 19.92 9.98 21.38 19.95 12.14 22.18 23.34
dbpedia 31.80 37.1 26.78 31.53 28.84 37.71 31.12 28.30 37.59 41.63
fever 65.13 68.6 53.42 56.28 52.35 71.11 42.06 51.75 63.97 74.06
fiqa 23.61 43.2 28.56 27.11 20.33 32.40 30.80 22.16 34.66 35.35
hotpotqa 63.30 52.2 52.37 19.64 44.75 47.05 24.32 42.12 48.51 65.29
nfcorpus 32.18 33.7 26.28 29.56 28.18 32.98 33.84 29.74 36.08 37.64
nq 30.55 41.7 37.65 34.43 29.55 43.14 38.25 30.37 46.97 48.30
quora 78.86 86.1 84.64 72.55 68.27 80.45 76.14 68.77 82.56 85.83
scidocs 14.90 21.8 10.39 18.51 11.57 17.59 20.59 13.25 19.10 18.82
scifact 67.89 72.3 66.36 52.68 58.48 65.71 63.12 61.53 70.34 73.58
trec-covid 59.47 61.8 63.34 59.52 54.59 69.25 67.64 63.00 76.85 80.29
touche 44.22 19.8 12.82 14.85 18.47 21.78 15.56 18.65 22.35 34.15
avg 43.70 44.61 41.38 35.87 34.48 44.13 38.08 35.88 45.88 50.60
@@
    Lowercase the input document text to align with the phrase "Make sure your word is in lowercase." in the prompt since this phrase skewed the sampling distribution towards lowercase tokens (a "sparser" distribution). We then utilize the NLTK toolkit Bird and Loper (2004) to extract all words in the document, filtering out standard English stopwords and punctuation.
@@
    Next, we use the LLM’s tokenizer to tokenize each extracted word and obtain their token IDs [^5] Note that many words may be split into sub-tokens, resulting in multiple token IDs, all of which are considered in the logits. We retain only the values corresponding to the obtained token IDs in the logits and set the rest of the dimensions to zero, thereby considering only tokens present in the documents, thus enabling exact term matching in retrieval.
@@
For dense retrieval, we directly use the hidden states as the embeddings of the documents. For indexing these embeddings, we simply normalize all the embeddings and add them into an Approximate Nearest search (ANN) vector index.
At query time, we process the queries exactly the same as the documents, with the only exception being that the term "passage" in the prompt is replaced with "query". The dense representation of the query is utilized for semantic search via the ANN index, while the sparse representation of the query is employed for exact term matching via the inverted index. Following previous work Wang et al. (2021b), we compute the final document scores by applying min-max normalization to both dense and sparse document scores. These normalized scores are then linearly interpolated with equal weights to produce the final document scores.
@@ 4 Experimental setup
Dataset and evaluation: We evaluate the document ranking effectiveness of both baseline methods and our proposed PromptReps using MSMARCO Nguyen et al. (2016) passage retrieval, TREC deep learning Craswell et al. (2020) and BEIR Thakur et al. (2021). These datasets encompass various IR tasks, providing a heterogeneous evaluation environment. For MSMARCO we report MRR@10 and for TREC deep learning and BIER we report nDCG@10 scores, the commonly employed evaluation measures for these datasets.
Baselines: We compare PromptReps with strong unsupervised first-stage retrievers including BM25, a classic term frequency-based sparse retrieval method, and E5-PTlarge Wang et al. (2024a), a state-of-the-art BERT-based dense embedding method trained on 1.3B carefully crafted unsupervised text pairs. LLM2Vec BehnamGhader et al. (2024), a Llama3-8B-Instruct LLM-based dense embedding method trained with bi-directional attention, masked next token prediction, and SimCSE Gao et al. (2021) on the Wikipedia corpus.
Implementation of PromptReps: PromptReps is implemented using four base LLMs: Mistral-7b-Instruct-v0.2666https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 Jiang et al. (2023a), Phi-3-mini-4k-instruct [^7] https://huggingface.co/microsoft/Phi-3-mini-4k-instruct Abdin et al. (2024), Llama3-8B-Instruct [^8] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct, and Llama3-70B-Instruct [^9] https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct AI@Meta (2024). Dense and sparse document and query encodings are implemented using the Huggingface Transformers library Wolf et al. (2020) and the Tevatron toolkit Gao et al. (2023b). The Faiss library Douze et al. (2024) is used to build the ANN index with cosine similarity as the embedding distance metric, and Pyserini Lin et al. (2021) is utilized to construct the inverted index for sparse retrieval. For the dense and sparse ranking hybrid, the Ranx library Bassani and Romelli (2022) is employed. In our experiments, we report dense only, sparse only, and the full hybrid results.
5 Results

### v3

@@ Abstract
Utilizing large language models (LLMs) for zero-shot document ranking is done in one of two ways: (1) prompt-based re-ranking methods, which require no further training but are only feasible for re-ranking a handful of candidate documents due to computational costs; and (2) unsupervised contrastive trained dense retrieval methods, which can retrieve relevant documents from the entire corpus but require a large amount of paired text data for contrastive training. In this paper, we propose PromptReps, which combines the advantages of both categories: no need for training and the ability to retrieve from the whole corpus. Our method only requires prompts to guide an LLM to generate query and document representations for effective document retrieval. Specifically, we prompt the LLMs to represent a given text using a single word, and then use the last token’s hidden states and the corresponding logits associated with the prediction of the next token to construct a hybrid document retrieval system. The retrieval system harnesses both dense text embedding and sparse bag-of-words representations given by the LLM. Our experimental evaluation on the MSMARCO, TREC deep learning and BEIR zero-shot document retrieval datasets illustrates that this simple prompt-based LLM retrieval method can achieve a similar or higher retrieval effectiveness than state-of-the-art LLM embedding methods that are trained with large amounts of unsupervised data, especially when using a larger LLM.
@@
Shengyao Zhuang1,2, Xueguang Ma3, Bevan Koopman1,2, Jimmy Lin3, Guido Zuccon2 1CSIRO, 2 The University of Queensland, 3 University of Waterloo
Figure 1: Overview of PromptReps. LLMs are prompted to simultaneously generate dense and sparse representations, then used to build search indexes.
@@ 1 Introduction
One line of work focuses on directly prompting LLMs to infer document relevance to a given query Sachan et al. (2022); Zhuang et al. (2023); Ma et al. (2023); Sun et al. (2023); Pradeep et al. (2023); Zhuang et al. (2024); Qin et al. (2024). For instance, RankGPT Sun et al. (2023) casts document re-ranking as a permutation generation task, prompting LLMs to generate re-ordered document identifiers according to the document’s relevance to the query. These methods leverage LLMs for document ranking in a complete zero-shot setting where no further training is required. However, these methods can only serve as a second-stage re-ranker on a handful of candidate documents. This is because each prompt requires one full LLM inference: for example, in the case of a corpus with 1M documents, a pointwise approach would require the construction of 1M prompts and thus the execution of 1M (costly) LLM inferences – making it unfeasible for an online search engine.
@@
Code for fully reproducing our results is available at https://github.com/ielab/PromptReps.
@@ 2.1 Supervised neural retrievers
Neural retrievers based on the bi-encoder architecture bring significant improvements over traditional best-match retrievers such as BM25. Dense retrievers such as DPR Karpukhin et al. (2020) and ANCE Xiong et al. (2021) are based on transformer language models and encode text into low-dimensional vectors, conducting search with nearest neighbor search. On the other hand, sparse neural retrievers such as DeepImpact Mallia et al. (2021), uniCOIL Lin and Ma (2021), TILDE Zhuang and Zuccon (2021c, b), and SPLADE Formal et al. (2021), also based on transformer language models, encode text into high-dimensional sparse vectors as bag-of-words representations, conducting search with inverted index. Recent works also explored fine-tuning generative LLMs as dense retrievers such as RepLLaMA Ma et al. (2024) and LLaRA Liao et al. (2024). A hybrid neural retrieval system refers to a system that combines the rankings provided by both dense and sparse retrievers, often resulting in an enhanced final ranking Lin and Ma (2021); Wang et al. (2021).
@@
All these retrievers are trained with supervised relevance judgment data (e.g., MSMARCO Bajaj et al. (2018)) using contrastive learning. Our work instead focuses on building a hybrid neural retrieval system with zero-shot dense and sparse document representations without supervised contrastive learning and based on generative LLMs. This capability has two implications: (1) no contrastive training is required, which is expensive when applied to LLMs with several billions parameters, and (2) no human-labelled training data is required, which may be laborious and expensive to obtain. With regards to the first point, Wang et al. (2024b) reported that the training of E5-mistrail (7B parameters) took about 18 hours on 32 V100 GPUs, for an approximate cost of USD \$2,300, [^1] Based on 4 On-Demand p3dn.24xlarge instances, June 2024. emissions of $\approx$5.6 kgCO2e and consumption of $\approx$37.7 L of water for the associated cooling activities. [^2] Emissions and water consumption estimates obtained using the frameworks of Scells et al. (2022); Zuccon et al. (2023). Scaling this training to more and larger LLMs, and more data, will consequently further increase costs. Our proposed method does not incur these additional contrastive pre-training costs. With regards to the second point, dense retrievers have shown to have poor generalisability when applied to data out-of-domain or out-of-task compared to the data used for contrastive training Thakur et al. (2021); Zhuang and Zuccon (2021a, 2022); Ren et al. (2023); Lin et al. (2023); Lupart et al. (2022). In presence of shift in data between training and deployment, retrieval losses can be significant: dense retrieval effectiveness can plummet far below that of best-match models like BM25 Khramtsova et al. (2023, 2024). The acquisition of in-domain/in-task training data can be costly, laborious and impractical/impossible especially in domain-specific applications when dealing with sensitive, private data.
@@ 2.2 Unsupervised neural retrievers
There have also been attempts at training effective neural retrievers without relying on human relevance judgments. Methods such as Contriever Izacard et al. (2022) and E5 Wang et al. (2024a), train a dense retriever with large-scale pseudo query-document pairs to build unsupervised training data. LLMs have also been adapted as unsupervised text embedding models for first-stage document retrieval. For instance, HyDE Gao et al. (2023a) enhances query representations for an unsupervised retriever by replacing the original query with LLM-generated hypothetical documents.
More recent work has focused on directly converting generative LLMs into a text-embedding model with unsupervised contrastive pre-training. Methods like E5-Mistral-Inst Wang et al. (2024b) and Gecko Lee et al. (2024) use large-scale weakly supervised paired text data or LLM-generated query-document pair data to perform contrastive training on top of LLMs. LLM2Vec BehnamGhader et al. (2024), on the other hand, conducts further masked next token prediction pre-training with bidirectional attention, and SimCSE Gao et al. (2021) trains on raw text data to transform LLMs into text encoders. Although no labeled data is used, these methods require synthetic or unsupervised paired text data to perform contrastive pre-training (thus still experiencing training costs in terms of computations; and further computational costs may be associated with the generation of synthetic training data). Our method instead relies solely on prompt engineering to transform LLM into a robust text encoder for document retrieval without any extra training.
@@ 2.3 Prompting LLMs for document ranking
Inspired by the prompt-following capacity of LLMs, recent studies have explored prompting LLMs for document re-ranking. For instance, UPR Sachan et al. (2022) ranks documents pointwise by prompting the LLM to generate a relevant query for a given document and rank documents based on the likelihood of generating the query. RankGPT Sun et al. (2023) and LRL Ma et al. (2023) propose to re-rank a list of documents at once and generate permutations for the reordered list. Pairwise Qin et al. (2024) and Setwise Zhuang et al. (2024) prompting methods have also been explored to improve effectiveness and efficiency in the LLM re-ranking pipeline. These methods are only feasible for re-ranking a handful of candidate documents, thus limited to second-stage document re-ranking. In contrast, our approach utilizes prompts to construct the first-stage retrievers.
2.4 Prompting LLMs for sentence embedding
@@ 3 PromptReps
Table 1: nDCG@10 scores of BEIR 13 publicly available datasets.
Sup Contrastive training Unsup Contrastive training PromptReps (ours)
LLM - BERT110M BERT110M BERT330M Llama3-8B-I Llama3-8B-I Llama3-70B-I
[…]
arguana 39.70 52.1 46.9 44.4 51.73 29.70 22.85 33.32 31.65 24.66 35.27
climatefever 16.51 22.8 22.7 15.7 23.58 19.92 9.98 21.38 19.95 12.14 22.18
dbpedia 31.80 44.2 41.7 37.1 26.78 31.53 28.84 37.71 31.12 28.30 37.59
fever 65.13 79.6 78.1 68.6 53.42 56.28 52.35 71.11 42.06 51.75 63.97
fiqa 23.61 35.1 35.6 43.3 28.56 27.11 20.33 32.40 30.80 22.16 34.66
hotpotqa 63.30 68.6 66.2 52.2 52.37 19.64 44.75 47.05 24.32 42.12 48.51
nfcorpus 32.18 34.5 33.9 33.7 26.28 29.56 28.18 32.98 33.84 29.74 36.08
nq 30.55 54.4 53.7 41.7 37.65 34.43 29.55 43.14 38.25 30.37 46.97
quora 78.86 81.4 87.5 86.1 84.64 81.77 70.35 84.24 81.18 67.69 83.70
scidocs 14.90 15.9 15.9 21.9 10.39 18.51 11.57 17.59 20.59 13.25 19.10
scifact 67.89 69.9 67.9 72.3 66.36 52.68 58.48 65.71 63.12 61.53 70.34
trec-covid 59.47 71.1 75.9 62.1 63.34 59.52 54.59 69.25 67.64 63.00 76.85
touche 44.22 24.4 26.3 19.8 12.82 14.85 18.47 21.65 15.56 18.65 22.35
avg 43.70 50.3 50.2 46.06 41.38 36.58 34.64 44.43 38.47 35.80 45.97
@@
    Lowercase the input document text to align with the phrase “Make sure your word is in lowercase.” in the prompt since this phrase skewed the sampling distribution towards lowercase tokens (a “sparser” distribution). We then utilize the NLTK toolkit Bird and Loper (2004) to extract all words in the document, filtering out standard English stopwords and punctuation.
@@
    Next, we use the LLM’s tokenizer to tokenize each extracted word and obtain their token IDs. [^4] Note that many words may be split into sub-tokens, resulting in multiple token IDs, all of which are considered in the logits We retain only the values corresponding to the obtained token IDs in the logits and set the rest of the dimensions to zero, thereby considering only tokens present in the documents, thus enabling exact term matching in retrieval.
@@
For dense retrieval, we directly use the hidden states as the embeddings of the documents. For indexing these embeddings, we simply normalize all the embeddings and add them into an Approximate Nearest search (ANN) vector index. In Appendix A, we provide example Python code of generating dense and sparse representations with PromptReps.
At query time, we process the queries exactly the same as the documents, with the only exception being that the term “passage” in the prompt is replaced with “query”. [^5] The only exception in our experiments is the Quora dataset, which is a duplicate query search task. Therefore, we use the query prompt for both queries and documents. The dense representation of the query is utilized for semantic search via the ANN index, while the sparse representation of the query is employed for exact term matching via the inverted index. Following previous work Wang et al. (2021), we compute the final document scores by applying min-max normalization to both dense and sparse document scores. These normalized scores are then linearly interpolated with equal weights to produce the final document scores. We do not explicitly tune the weight because our setting is zero-shot retrieval, and we wanted to maintain the “zero-shot” nature of our approach. Nevertheless, in Appendix B, we explore the impact of the different weight settings.
@@ 4 Experimental Setup
Dataset and evaluation: We evaluate the document ranking effectiveness of both baseline methods and our proposed PromptReps using MSMARCO Bajaj et al. (2018) passage retrieval, TREC deep learning Craswell et al. (2020) and BEIR Thakur et al. (2021). These datasets encompass various IR tasks, providing a heterogeneous evaluation environment. For MSMARCO we report MRR@10 and for TREC deep learning and BIER we report nDCG@10 scores, the commonly employed evaluation measure for these datasets.
[…]
Implementation of PromptReps: PromptReps is implemented using four base LLMs: Mistral-7b-Instruct-v0.2666https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 Jiang et al. (2023a), Phi-3-mini-4k-instruct [^7] https://huggingface.co/microsoft/Phi-3-mini-4k-instruct Abdin et al. (2024), Llama3-8B-Instruct, [^8] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct and Llama3-70B-Instruct [^9] https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct AI@Meta (2024). Dense and sparse document and query encodings are implemented using the Huggingface Transformers library Wolf et al. (2020) and the Tevatron toolkit Gao et al. (2023b). The Faiss library Douze et al. (2024) is used to build the ANN index with cosine similarity as the embedding distance metric. We simply use brute force search for ANN (IndexFlatIP in Faiss) for a fair comparison with the baselines. For sparse retrieval, Pyserini Lin et al. (2021) is utilized to construct the inverted index. For the dense and sparse ranking hybrid, the Ranx library Bassani and Romelli (2022) is employed. In our experiments, we report dense only, sparse only, and the full hybrid results.
5 Zero-shot Results